OpenStudy (anonymous):

5/2 square root 6 minus 1 please help

OpenStudy (anonymous):

5\[\frac{ 5 }{ 2\sqrt{6}-1 }\]

OpenStudy (anonymous):

1.2823868446883383035640365542199527791156293394159435

OpenStudy (anonymous):

hehehehehe

OpenStudy (anonymous):

Are you wanting it to some simplified form? I don't understand what you are asking? \[\frac{ 5 }{ 2\sqrt(6)-1 }=\frac{ 10\sqrt(6)+1 }{ 23 }\]

OpenStudy (anonymous):

Whoops the numerator should be 10*sqrt(6)+5

OpenStudy (anonymous):

well im trying the decimal answer

OpenStudy (anonymous):

im not really sure how to put in on the calculator